OpenStudy (anonymous):

What is the slope of the line between (3, -4) and (-2, 1)?

OpenStudy (anonymous):

\[\frac{ y_2 - y_1 }{ x_2 - x_1 } = m = slope\]

OpenStudy (anonymous):

is the answer 2? @DemolisionWolf

OpenStudy (anonymous):

no..ans is -1... put values here in eq... 1-(-4)/-3-2=5/-5=-1